### v3

# Filtrated Kinematic Connectivity Analysis for Lower‑limb Joint Effective Age Evaluation

Abstract
To understand and communicate the risk of chronic lower-limb joint diseases associated with aging, it is crucial to investigate the relationship between age and gait dynamics, particularly through angular kinematics. One key challenge is that angular kinematic trajectories are highly interconnected, and the structures of the interconnections vary across different components. Neglecting the interconnections and the variability in the connectivity structures impairs the understanding of age‑associated gait coordination. To this end, we develop a novel kinematic connectivity analysis framework, grounded in multiple functional regression, to evaluate lower-limb joint effective age and uncover age-related kinematic features. The proposed approach is built upon the concept of filtration, a widely used tool in network analysis and topological data analysis for multi-resolution exploration. Specifically, we develop a forest-structured covariate grouping framework in which different kinematic trajectories are aggregated hierarchically to capture both (partially) shared and idiosyncratic motion signatures which are strongly associated with aging. We also develop a novel filtrated functional partial least squares approach for model estimation and feature extraction. Compared to existing approaches, our proposed approach demonstrates superior predictive power while providing novel insights into the coordinated evolution of angular kinematics during aging. In addition, the proposed framework is broadly applicable and can be readily extended in other scientific domains.
Keywords: Angular kinematics, Effective age, Filtrated grouping, Multiple functional regression, Partial least squares.
1 Introduction
1.1 Backgrounds, Data Description, and Objectives
Gait analysis (GA) is a widely utilized method for examining and understanding angular kinematics, which involves studying the rotational movements of lower-limb joints during activities such as walking. This analysis provides valuable insights into joint function and coordination, aiding in the assessment of movement patterns and the identification of potential abnormalities (see e.g., Kay et al. [20], Lofterød et al. [25] and Rao et al. [31]). As individuals age, changes in joint flexibility, muscle strength, and neuromuscular control can alter these angular movements, leading to decreased mobility and increased fall risk. The study in this paper is driven by a growing scientific interest in the intersection of chronic risk communication and gait analysis.
An emerging concept, referred to as “effective age” has garnered increasing attention (see e.g., Anderson and Loeser [2], Loeser [24], Li et al. [23], and Spiegelhalter [35]). Effective age represents the average age level of healthy individuals who share the same risk profile as a person being evaluated. In many cases, a person’s health status is more accurately reflected by their effective age rather than their chronological age, as it accounts for the impact of chronic disease risks and overall well-being. Evaluating effective age through gait analysis is a vital component in the proactive management of joint health, facilitating early detection of degeneration, customization of rehabilitation strategies, and the promotion of sustained mobility and independence.
Importantly, coordination among different lower-limb angular movement is pivotal for understanding human movement patterns and their underlying biomechanics. Existing research underscores the importance of inter-joint connectivity in angular kinematics as essential for maintaining balance, stability, and forward progression (see e.g., Ren et al. [32] and Dussault-Picard et al. [10]). Such coordination is crucial for evaluating effective age. As revealed in our analysis, it gives rise to homogeneous age‑associated patterns across different kinematic trajectories, a phenomenon we refer to as kinematic connectivity. Studying homogeneous age‐related associations across kinematic trajectories helps uncover systemic coordinated changes in gait that are more informative and clinically relevant than examining each kinematic trajectory in isolation. It reflects comprehensive aging mechanisms rather than noisy joint‐specific effects. In addition, the connectivity enables simplified model structures and improving evaluation performance.
There are several challenges to solve: 1) the kinematic connectivity structure underlying the association between age and kinematic trajectories is unknown, and should be appropriately identified to achieve accurate and interpretable effective age evaluations. Imposing a fixed connectivity structure as prior knowledge often leads to model misspecification, and such model assumptions are difficult to justify; 2) Since kinematic trajectories are functional data, the kinematic connectivity may change across different dimensions, the developed analytical framework should be tailored to tackle this variability; 3) While kinematic connectivity exists, the idiosyncratic features pertaining to each individual kinematic trajectory also play a crucial role in reflecting lower-limb joint aging. Therefore, distinguishing the shared (homogeneous) components from the trajectory-specific (idiosyncratic) components is essential. All these challenges motivate us to develop a new analytical framework, capable of uncovering kinematic connectivity with complex structures.
To summarize, this study aims to address the following objectives:
  • 1.
    We aim to develop an effective age assessment framework characterizing joint aging mechanisms using kinematic features constructed from angular kinematic trajectories of pelvic, hip, knee, ankle and foot.
  • 2.
    We aim to detect the consistent age‑associated patterns among the angular kinematic trajectories, thereby offering a system‑level view of musculoskeletal aging, and to identify both (partially) homogeneous and idiosyncratic kinematic features that significantly inform effective age evaluation.
Here we analyze a dataset generated from the experiment conducted by Fukuchi et al. [14]. The study involves 42 healthy volunteers aged between 21 and 84, all of whom had no lower-extremity injuries in the six months prior to participation and no orthopedic or neurological conditions affecting their gait. Since the number of participants is limited, it is crucial to avoid building models with high complexity. Studying the association between chronological age and gait angular kinematics in healthy individuals is essential, as it serves as a baseline for assessing new subjects. Although chronological age and effective age do not always align in healthy individuals, particularly in those engaging in regular exercise that can reduce biological age by several years, the phenomenon is not universal and should not be assumed to occur in most healthy people. Data were collected using standard gait-analysis procedures in a 10×12-m room at the Laboratory of Biomechanics and Motor Control at the Federal University of ABC, Brazil. This facility is equipped with a motion-capture system that had 12 cameras, five force platforms embedded in the floor, and a dual-belt instrumented treadmill. Our study focuses on treadmill walking data, as they offer greater consistency and are not influenced by terrain variations. The angular kinematics data were recorded at various gait speeds, categorized into eight levels (S1 — S8) ranging from the slowest to the fastest speeds. The ensemble averages of angular kinematics across different speed levels are shown in Figure 1.
Figure 1: Angular kinematics (right-hand side) during treadmill walking averaged over different levels of gait speed. S5 represents the comfortable walking speed.
1.2 Functional Linear Regression and Covariate Grouping
Multivariate functional data analysis (FDA) offers a powerful toolbox for handling multiple complex functional objects. Its primary strength lies in addressing the dual challenges of high dimensionality inherent in individual functional elements and the intricate connectivity among them. Recent advancements in this field have enabled researchers to tackle diverse challenges across various scientific domains, including neuroimaging analysis (e.g., Wang et al. [36] and Jiao et al. [18]), public health (e.g., Zhou et al. [41]), and gait analysis (e.g., Cheng et al. [4], Dannenmaier et al. [8], and Cao et al. [3]). These developments underscore the potential of multivariate FDA in addressing complex high-dimensional data. Among its various tools, functional linear regression (see Ramsay and Silverman [30] and Morris [27] for overviews) stands out as a well-established class of linear models. In this article, we focus on the following model with a scalar response $y_{n}\in\mathbb{R}$ and multiple functional covariates $\{X_{jn}(t)\in L^{2}[0,1]\colon j=1,\ldots,p\}$,
$y_{n}=\beta_{0}+\sum_{j=1}^{p}\langle X_{jn},\beta_{j}\rangle+\epsilon_{n},\\ \ \mbox{E}\epsilon_{n}=0,\ \mbox{Var}(\epsilon_{n})=\sigma^{2},$ (1-1)
where the coefficient functions $\{\beta_{j}(t)$, $j=1,\ldots,p\}\in(L^{2}[0,1])^{p}$, and $\{(X_{jn}(t),\epsilon_{n})\colon j\geq 1\}$ are independent across $n$. More research works on functional linear models include, but are not limited to, Chiou et al. [5, 6, 7], Müller and Stadtmüller [28], Müller and Yao [29], Yao et al. [38], Ferraty et al. [13], Jiao and Ombao [19], Jiao et al. [16].
As discussed previously, besides evaluating joint effective age, an important task is to account for the connectivity among different covariates (i.e., the angular kinematic trajectories of various joints). Such connectivity leads to covariates being associated with the response variable (i.e., chronological age) in a homogeneous manner, which is crucial to investigate the homogeneous angular kinematic evolution in relation to age. A reasonable strategy to tackle the connectivity is to group different kinematic trajectories based on the similarity of their association with age, and then extract the homogeneous association patterns based on the grouping structure. This first step is known as grouping pursuit (see Shen and Huang [34], Ke et al. [21], Ma and Huang [26], She et al. [33], and Jiao and Chan [17]). However, ordinary grouping pursuit methods face a fundamental limitation: they assume a uniform grouping structure and thus cannot accommodate variability in kinematic connectivity across different dimensions.
1.3 Contributions and Novelties
Methodologically, we develop a filtration‑based grouping algorithm for model (1-1), termed “filtrated grouping”. This idea is inspired by filtration, a widely used technique in graph analysis and topological data analysis (see, e.g., Edelsbrunner et al. [11], Lee et al. [22], and Hofer et al. [15]), which iteratively constructs a sequence of nested structures and serves as a fundamental building block for multi-resolution analysis. Rather than isolating the effects of individual covariates, our approach captures both shared (homogeneous) and distinct (idiosyncratic) association patterns across multiple levels of resolution. The underlying principle is that as homogeneous components are progressively extracted from different functional covariates, the remaining parts should increasingly exhibit idiosyncratic associations. The covariates are grouped iteratively in each layer of filtration, leading to dimension-specific grouping structures. This new analysis framework not only enhances model interpretability but also improves the accuracy of effective age evaluation. Filtration techniques have rarely been applied to functional data. To the best of our knowledge, the only related work is Jiao et al. [18], which employs filtration to construct a multi-resolution structure for partial common functional principal components (PCfPC), offering a more inclusive and efficient representation of PCfPCs. This paper is the first work to develop filtration techniques for multiple functional regression.
In the application, we develop an innovative tool to study the association between joint angular kinematics and chronological age, serving as a foundation for effective age evaluation and offering new insights into kinematic connectivity. To achieve this, we use multiple functional regression, where age serves as the response variable and angular kinematic curves are treated as functional predictors. By employing our new grouping algorithm, both shared and idiosyncratic kinematic features are effectively captured in a novel hierarchical manner. The variability in the kinematic connectivity structures is effectively addressed, and our analysis highlights the key kinematic features relevant to the assessment of effective age. The findings deepen our understanding of the coordination in age‑related evolution of angular gait kinematics and provide a valuable tool for assessing chronic lower‑limb joint risks.
The rest of the article is organized as follows. We present the filtrated grouping structure and detail the model estimation and feature extraction procedures in Section 2. In Section 4, we apply our methods to analyze the association between chronological age and angular kinematics data. We discuss the results in Section 5. Simulation results and additional algorithms are presented in the supplementary materials.
2 Methodology
2.1 Preliminaries
Notationally, let $X\in L^{q}_{H}=L_{H}^{q}(\Omega,\mathcal{A},\mathcal{P})$ indicate that, for some $q>0$, a $H$-valued function $X$ satisfies $E\|X(t)\|^{q}<\infty$. In what follows, we assume that there are one scalar response $y_{n}$ representing chronological age and $p$ functional covariates $\{X_{jn}(t)\in L^{2}_{H}\colon j=1,\ldots,p\}$ representing different kinematic trajectories, where $H=L^{2}[0,1]$ is equipped with the inner product $\langle x,y\rangle=\int_{0}^{1}x(t)y(t)dt$ and the norm $\|x\|^{2}=\int_{0}^{1}x^{2}(t)dt$. It is assumed that $\{y_{n},X_{jn}(t)\colon j=1,\ldots,p\}$ are independent across $n=1,\ldots,N$, where $N$ is the sample size. By Mercer’s theorem, we have the following functional principal component decompositions $\mathcal{C}_{j}(\cdot)=\sum_{d\geq 1}\theta_{jd}\langle\nu_{jd},\cdot\rangle\nu_{jd}(t),$ where $\mathcal{C}_{j}(x)(t)=\int C_{j}(t,s)x(s)ds$ for $x\in L^{2}[0,1]$, $\mu_{j}(t)=\mbox{E}X_{jn}(t)$, and $C_{j}(t,s)=\mbox{E}\{(X_{jn}-\mu_{j})(t)(X_{jn}-\mu_{j})(s)\}$. We assume that all the coefficient functions in (1-1) belong to the space spanned by $\{\nu_{jd}(t)\colon d\geq 1,\ j\geq 1\}$.
Suppose that, for some sets of basis functions $\{\nu_{jd}(t)\in H\colon d\geq 1\}$ and $\{u_{jd}(t)\in H\colon d\geq 1\}$, the functional covariates and functional coefficients admit the following basis representation
$X_{jn}(t)=\sum_{d\geq 1}\xi_{jn,d}\nu_{jd}(t),\ \beta_{j}(t)=\sum_{d,d^{\prime}\geq 1}b_{jd}M_{j,dd^{\prime}}u_{jd^{\prime}}(t),$
in which the basis functions $\{\nu_{jd}(t)\colon d\geq 1\}$ and $\{u_{jd}(t)\colon d\geq 1\}$ can either be orthonormal or not. The infinite matrix $\{M_{j,dd^{\prime}}\colon d,d^{\prime}\geq 1\}$ satisfies the condition $\sum_{d^{\prime}\geq 1}M_{j,d_{1}d^{\prime}}(\int\nu_{jd_{2}}u_{jd^{\prime}})=\bm{1}_{\{d_{1}=d_{2}\}}$. Then model (1-1) can be equivalently rewritten in the following form
$y_{n}=\beta_{0}+\sum_{d=1}^{\infty}\sum_{j=1}^{p}\xi_{jn,d}b_{jd}+\epsilon_{n},\\ \ \mbox{E}\epsilon_{n}=0,\ \mbox{Var}(\epsilon_{n})=\sigma^{2}.$ (2-2)
For illustrating the new filtrated grouping structure, we focus on the equivalent form (2-2), as the representation is dimension-specific (layer-specific), enabling a more precise characterization of grouping structures across different dimensions.
2.2 Filtrated Grouping Structure
In this study, we model kinematic connectivity using a hierarchical grouping strategy. To begin, we introduce the ordinary grouped regression model as a foundation. Given a grouping structure $G=(\mathcal{K}_{1},\ldots,\mathcal{K}_{m})$ satisfying $\mathcal{K}_{i}\cap\mathcal{K}_{i^{\prime}}=\emptyset$ for $i\neq i^{\prime}$ and $\cup_{i=1}^{m}\mathcal{K}_{i}=\{1,\ldots,p\}$, the ordinary grouped model based on coefficient homogeneity is given as follows
$y_{n}=\beta_{0}+\sum_{i=1}^{m}\langle Z_{in},\alpha_{i}\rangle+\epsilon_{n},\\ \ \mbox{E}\epsilon_{n}=0,\ \mbox{Var}(\epsilon_{n})=\sigma^{2},$ (2-3)
where $Z_{in}(t)=\sum_{j\in\mathcal{K}_{i}}X_{jn}(t)$ and $\alpha_{i}(t)$ is the common coefficient function of group $\mathcal{K}_{i}$, capturing the homogeneous association shared by the covariates within the group. The ordinary model (1-1) cannot account for the homogeneity of functional coefficients and the estimation burden is high even for moderate $p$. Although the grouped model (2-3) alleviates the estimation burden and improve model interpretability, it fails to explain the data mechanism when the grouping structures are different across different dimensions. Indeed, as demonstrated in Section 4, both common and idiosyncratic components of different angular kinematic trajectories play important roles in effective age evaluation, and imposing a uniform grouping structure results in suboptimal model performance. This motivates the development of the filtrated grouping procedure.
Building upon the equivalent model representation (2-2), we develop the following filtrated grouped model
$y_{n}=\beta_{0}+\sum_{d=1}^{\infty}\sum_{i=1}^{m_{d}}\left\{\sum_{j\in\mathcal{K}_{d,i}}\xi_{jn,d}\right\}a_{d,i}+\epsilon_{n},\ \mbox{E}\epsilon_{n}=0,\ \mbox{Var}(\epsilon_{n})=\sigma^{2},$ (2-4)
where $\mathcal{K}_{d,i}$ denotes the $i$-th group in the $d$-th dimension/filtration layer and $m_{d}$ is the number of groups in the $d$-th dimension. Compared to the ordinary grouped model (2-3), the filtrated grouped model (2-4) accommodates grouping variability across different dimensions. The homogeneous components are defined as follows
$\zeta_{nd}^{(i)}\overset{\vartriangle}{=}\sum_{j\in\mathcal{K}_{d,i}}\xi_{jn,d},\ i\geq 1,\ d\geq 1.$
These components capture the features shared among covariates within each group, which are expected to have strong associations with the response. In particular, when $\mathcal{K}_{d,i}$ contains only a single covariate, $\zeta_{nd}^{(i)}$ captures the idiosyncratic feature of the covariate. In addition to obtaining these homogeneous components, we also need to estimate the associated coefficient scores $\{a_{d,i}\colon i\geq 1,\ d\geq 1\}$.
The homogeneous components are extracted in a hierarchical manner, resulting in group indices $\{\mathcal{K}_{d,i}\colon i\geq 1,d\geq 1\}$ that follow a forest structure (see Figure 2). In the first layer, we identify the most homogeneous covariate components and group them accordingly to obtain $\{\zeta_{n1}^{(i)}\colon i\geq 1\}$ and estimate $\{a_{1,i}\colon i\geq 1\}$. After filtering out these homogeneous components from each covariate, we divide $\{\mathcal{K}_{1,i}\colon i\geq 1\}$ into $\{\mathcal{K}_{2,i}\colon i\geq 1\}$, which are either nested within or identical to the groups in the previous layer (i.e., for any $i\geq 1$, $\mathcal{K}_{2,i}\subseteq\mathcal{K}_{1,i^{\prime}}$ for some $i^{\prime}\geq 1$), and then obtain the homogeneous components in the new layer (i.e., $\{\zeta_{n2}^{(i)}\colon i\geq 1\}$) and estimate the associated coefficient scores (i.e., $\{a_{2,i}\colon i\geq 1\}$). The splitting procedure continues until a predefined stopping criterion is met. In the deeper layers of the filtration, higher-resolution homogeneous components, constructed from fewer covariates, are progressively extracted.
Figure 2: Hierarchical forest structure of coefficient homogeneity. Here, $L_{d}$ signifies the $d$-th layer of filtration. Different groups are represented by different colored solid circles. Each group is either nested within or identical to its “parent” group (i.e., the group from which it originates through a split) in the previous layer.
In the remainder of the section, we address the following problems: Given the forest structure, how to extract the homogeneous components and estimate the coefficient scores? How to determine a reasonable forest structure for filtrated covariate grouping? In Section 2.3, we address the first problem by developing a filtrated functional partial least squares approach. In Section 2.4, we construct the forest structure based on the grouping path derived from the grouping pursuit problem.
2.3 Filtrated Functional Partial Least Squares (filt-fPLS)
To extract the homogeneous components and estimate the coefficient scores, we develope a filtrated functional partial least squares method. To facilitate understanding of the proposed filt-fPLS algorithm, we begin with a simple illustrative case. In the ordinary grouped model (2-3), if $X_{1n}(t),\ldots,X_{pn}(t)\in\mathcal{K}_{1}$ (i.e., all covariates belong to a single group $\mathcal{K}_{1}$), then we can estimate the model using the classical functional partial least squares approach (see Delaigle and Hall [9]), where $\sum_{j\in\mathcal{K}_{1}}X_{jn}(t)$ is treated as the covariate. The first step is to maximize $\mbox{cov}(y_{n},\sum_{j\in\mathcal{K}_{1}}\langle X_{jn},\psi\rangle)$ over $\psi(t)$, subject to the constraint $\|\psi\|=1$. The resulting feature $\sum_{j\in\mathcal{K}_{1}}\langle X_{jn},\psi\rangle$ is used to explain the response. This is followed by a projection step, and the two steps are then repeated iteratively. In the filtrated grouped model (2-4), however, the grouping structure varies across dimensions. As a result, the maximization and projection steps are iteratively performed under different grouping configurations. The core principle of filt-fPLS is that the extracted homogeneous components should exhibit strong explanatory power for the response variable. We detail the algorithmic procedure below.
When $d=1$, we first find the filt-fPLS basis for each group as follows,
$\psi_{1,i}(t)\overset{\vartriangle}{=}\arg\max_{\|\psi\|=1}\mbox{cov}\left\{y_{n},\left\langle\sum_{j\in\mathcal{K}_{1,i}}X_{jn},\psi\right\rangle\right\}=\frac{\mbox{cov}\{y_{n},\sum_{j\in\mathcal{K}_{1,i}}X_{jn}\}}{\|\mbox{cov}\{y_{n},\sum_{j\in\mathcal{K}_{1,i}}X_{jn}\}\|}(t),$
where $i=1,\ldots,m_{1}$. The $m_{1}$ first-layer homogeneous components are obtained as $\zeta_{n1}^{(i)}=\sum_{j\in\mathcal{K}_{1,i}}\int X_{jn}\psi_{1,i}$ for $i=1,\ldots,m_{1}$, and then estimate the associated coefficient scores:
$\hat{a}_{1,i}=\arg\min_{\{a_{1,i}\colon i=1,\ldots,m_{1}\}}\mbox{E}\left\{y_{n}-\sum_{i=1}^{m_{1}}a_{1,i}\zeta_{n1}^{(i)}\right\}^{2}.$
For $d\geq 2$, given $\{\psi_{d-1,i}(t)\colon i=1,\ldots,m_{d-1}\}$, the filt-fPLS basis for group $\mathcal{K}_{d,i}$ is
$\psi_{d,i}(t)\overset{\vartriangle}{=}\arg\max_{\|\psi\|=1}\mbox{cov}\left\{y_{n}^{[d]},\left\langle\sum_{j\in\mathcal{K}_{d,i}}X^{[d]}_{jn},\psi\right\rangle\right\}=\frac{\mbox{cov}\{y_{n}^{[d]},\sum_{j\in\mathcal{K}_{d,i}}X^{[d]}_{jn}\}}{\|\mbox{cov}\{y_{n}^{[d]},\sum_{j\in\mathcal{K}_{d,i}}X^{[d]}_{jn}\}\|}(t).$
Let $\zeta_{nd}^{(i)}=\sum_{j\in\mathcal{K}_{d,i}}\int X^{[d]}_{jn}\psi_{d,i}$, then obtain the associated coefficient scores as follows
$\hat{a}_{d,i}=\arg\min\limits_{\{a_{d,i}\colon i=1,\ldots,m_{d}\}}\mbox{E}\left\{y^{[d]}_{n}-\sum_{i=1}^{m_{d}}a_{d,i}\zeta_{nd}^{(i)}\right\}^{2},$
where
$X^{[d]}_{jn}(t)$ $=X^{[d-1]}_{jn}(t)-\sum_{i=1}^{m_{d-1}}\zeta_{n,d-1}^{(i)}\phi_{d-1,i}(t),$
$\phi_{d,i}(t)$ $=\arg\min_{\{\delta_{i}(t)\colon i\geq 1\}}\mbox{E}\left\|X^{[d]}_{jn}(t)-\sum_{i=1}^{m_{d}}\zeta_{nd}^{(i)}\delta_{i}(t)\right\|^{2},$
$y_{n}^{[d]}$ $=y_{n}^{[d-1]}-\sum_{i=1}^{m_{d-1}}\zeta_{n,d-1}^{(i)}\hat{a}_{d-1,i}.$
Particularly, $X_{jn}^{[1]}(t)=X_{jn}(t)$, $y_{n}^{[1]}=y_{n}$.
In each layer, the filt‑fPLS basis functions and the corresponding coefficient scores jointly account for homogeneous association patterns of covariates in the same group. Note that under forest structure, the homogeneous components across different layers are orthogonal (i.e., $\sum_{n\geq 1}\zeta_{nd}^{(i)}\zeta_{nd^{\prime}}^{(i^{\prime})}=0$ for $d\neq d^{\prime}$ and $i,i^{\prime}\geq 1$). Consequently, the homogeneous components identified at different layers capture distinct information, allowing the filtrated grouped model (2-4) to represent multi-resolution kinematic connectivity structure without redundancy and with clearer interpretability. In practical applications, the unknown covariances and expectations are replaced by their empirical counterparts, as discussed in Algorithm 1. In the next section, we discuss how to identify the forest structure.
Algorithm 1 Filtrated functional partial least squares algorithm
1:Set $d=1$, $X_{jn}^{[1]}(t)=X_{jn}(t)$, and $y_{n}^{[1]}=y_{n}$.
2:while the ending criterion (discussed in Section 2.4.3) is not met do
3:  For $i=1,\ldots,m_{d}$, estimate $\psi_{d,i}(t)$ by
$\hat{\psi}_{d,i}(t)=\sum_{n=1}^{N}\sum_{j\in\mathcal{K}_{d,i}}y_{n}^{[d]}X_{jn}^{[d]}(t)\bigg{/}\left\|\sum_{n=1}^{N}\sum_{j\in\mathcal{K}_{d,i}}y_{n}^{[d]}X_{jn}^{[d]}\right\|.$
4:  Set $\zeta_{nd}^{(i)}=\sum_{j\in\mathcal{K}_{d,i}}\langle X_{jn}^{[d]},\hat{\psi}_{d,i}\rangle$ and estimate the coefficient scores by least squares
$\hat{a}_{d,i}$ $=\arg\min_{\{a_{d,i}\colon i=1,\ldots,m_{d}\}}\sum_{n=1}^{N}\left\{y_{n}^{[d]}-\sum_{i=1}^{m_{d}}a_{d,i}\zeta_{nd}^{(i)}\right\}^{2}.$
5:  Obtain the residuals
$X_{jn}^{[d+1]}(t)$ $=X_{jn}^{[d]}(t)-\sum_{i=1}^{m_{d}}\zeta_{nd}^{(i)}\hat{\phi}_{d,i}(t),$
$\hat{\phi}_{d,i}(t)$ $=\arg\min_{\{\delta_{i}(t)\colon i\geq 1\}}\sum_{n=1}^{N}\left\|X_{jn}^{[d]}(t)-\sum_{i=1}^{m_{d}}\zeta_{nd}^{(i)}\delta_{i}(t)\right\|^{2},$
$y_{n}^{[d+1]}$ $=y_{n}^{[d]}-\sum_{i=1}^{m_{d}}\hat{a}_{d,i}\zeta_{nd}^{(i)}.$
6:  $d=d+1$
7:end while
2.4 Forest Structure Construction
The construction of the forest structure is based on the grouping path $\mathcal{P}$ derived from the ordinary grouping pursuit problem. Specifically, this involves identifying $\mathcal{K}_{1},\ldots,\mathcal{K}_{m}$ in model (2-3) by solving a regularized optimization problem over a sequence of increasing penalty parameters (see e.g., Shen and Huang [34]), and $\mathcal{P}$ encompasses all distinct grouping structures identified across different penalty parameters. We apply filtration along the grouping path because covariates that share strong homogeneity in their coefficient functions tend to merge earlier, providing useful information on the hierarchical structures of kinematic connectivity. Based on the grouping path, we first find a sequence of candidate grouping structures from which the forest structure is constructed. For the candidate grouping structures, it is required that the groups in each layer are nested within those of the preceding layer, as any non-nested grouping structures would complicate the construction of the forest structure. One challenge is that the grouping structures along the grouping path may not be nested as $\lambda$ increases, due to the occurrence of group splitting. Thus, it is crucial to carefully select the candidate grouping structures from the grouping path. In the following, we discuss the identification of grouping path in Section 2.4.1, the selection of candidate grouping structures in Section 2.4.2, and the construction of the forest structure in Section 2.4.3.
2.4.1 Grouping Path
The grouping pursuit problem is formulated by optimizing the objective function:
$S(\bm{\beta}(t),\lambda)=\frac{1}{2}\sum_{n=1}^{N}\left(y_{n}-\sum_{j=1}^{p}\langle X_{jn},\beta_{j}\rangle\right)^{2}+\sum_{i<j}J_{\lambda}(\|\beta_{i}-\beta_{j}\|),$
where $\bm{\beta}(t)=\{\beta_{j}(t)\colon j=1,\ldots,p\}$, and $J_{\lambda}(\cdot)$ is some concave penalty, such as the SCAD penalty (see Fan and Li [12]) and the MCP (see Zhang [39]). As $\lambda$ increases, covariates tend to be clustered into fewer groups, although occasional group splits may still occur. The sequence of all grouping structures that emerge as the regularization parameter $\lambda$ increases defines the grouping path $\mathcal{P}$. Suppose that for some orthonormal basis functions $\{\nu_{d}(t)\colon d\geq 1\}$, the following basis representations hold $X_{jn}(t)=\sum_{d\geq 1}\xi_{jn,d}\nu_{d}(t)$ and $\beta_{j}(t)=\sum_{d\geq 1}b_{jd}\nu_{d}(t)$. Since $S(\bm{\beta}(t),\lambda)$ involves infinite-dimensional elements, we minimize the following truncated version over $\{b_{jd}\colon d=1,\ldots,D,\ j\geq 1\}$ with some sufficiently large $D\geq 1$,
$S_{D}(\bm{\beta}(t),\lambda)=\frac{1}{2}\sum_{n=1}^{N}\left(y_{n}-\sum_{j=1}^{p}\bm{\xi}^{\top}_{jn}\bm{B}_{j}\right)^{2}+\sum\limits_{i<j}J_{\lambda}(\|\bm{B}_{i}-\bm{B}_{j}\|),$ (2-5)
where $\bm{\xi}_{jn}=(\xi_{jn,1},\ldots,\xi_{jn,D})^{\top}$, $\bm{B}_{j}=(b_{j1},\ldots,b_{jD})^{\top}$. We use the ADMM algorithm to minimize (2-5). The details are included in the supplementary materials.
2.4.2 Candidate Grouping Structures
After obtaining the grouping path $\mathcal{P}$, we first construct candidate sets of grouping structures from the path, from which the forest structure is then constructed. Notationally, for two grouping structures $^{\prime}, GGG\subseteq G^{\prime}$ means that every group in $G$ is nested in some group in $G^{\prime}$. Denote a candidate set by $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$, where $\ell$ is the number of grouping structures in $\mathcal{G}$, and $G_{i}\in\mathcal{P}$ denotes the $i$-th layer’s grouping structure of the candidate set $\mathcal{G}$ satisfying $G_{\ell}\subseteq G_{\ell-1}\subseteq\cdots\subseteq G_{1}$.
Multiple candidate sets can be constructed from the grouping path, and only one of them is used for constructing the forest structure of model (2-4). The relationship $\mathcal{G}\subseteqq\mathcal{G}^{\prime}$ means that every grouping structure in $\mathcal{G}$ is also included in $\mathcal{G}^{\prime}$. To make the candidate set as inclusive as possible, we require that any two different candidate sets $\mathcal{G}=\{G_{1},\ldots,G_{\ell}\}$ and $\mathcal{G}^{\prime}=\{G^{\prime}_{1},\ldots,G^{\prime}_{\ell^{\prime}}\}$ satisfy that a) $\mathcal{G}\not\subseteqq\mathcal{G}^{\prime}$ and $\mathcal{G}^{\prime}\not\subseteqq\mathcal{G}$, b) ${G}_{1}={G}^{\prime}_{1}=\{1,2,\ldots,p\}$, and ${G}_{\ell}={G}^{\prime}_{\ell^{\prime}}=\{\{1\},\{2\},\ldots,\{p\}\}$. The selected candidate set should ensure sufficient diversity in grouping structures to support forest construction and achieve strong predictive performance in the resulting model. We now present the procedure for selecting the candidate set.
  • 1.
    Initial Construction: Construct all candidate sets that satisfy conditions (a) and (b) from the grouping path $\mathcal{P}$.
  • 2.
    Set Evaluation: For each candidate set, we construct the grouped model (2-3) based on all grouping structures in the set and estimate the prediction error using a resampling-based method (e.g., cross-validation, bootstrap).
  • 3.
    Final Selection: For each candidate set, compute the average prediction error of the models built across all the included grouping structures, and select the set that yields the lowest average prediction error.
This selection procedure ensures that the chosen candidate set represents a comprehensive forest hierarchy, including both the finest and the coarsest grouping. Meanwhile, the grouping structures in the selected candidate set retain strong explanatory power, thereby aligning with the covariance maximization principle central to filt-fPLS.
2.4.3 Iterative Selection of Grouping Structures
The forest structure is constructed from the selected candidate set $\widetilde{\mathcal{G}}=\{\widetilde{G}_{1},\ldots,\widetilde{G}_{\ell}\}$ through an iterative process. If the grouping structure selected for the current layer is $\widetilde{G}_{i}$ based on a predefined criterion, the process continues to the next layer by sequentially testing $\widetilde{G}_{i},\ldots,\widetilde{G}_{\ell}$. Particularly, we sequentially examine all the grouping structures in $\widetilde{\mathcal{G}}$ in the first layer of filtration. Regarding the predefined criterion, we adopt the generalized information criterion (GIC, see e.g., Zhang et al. [40]) to select the grouping structures across different layers. The format of GIC criterion is: GIC = measure of model fit ++ tuning parameter $\times$ measure of model complexity. In principle, if the covariates are separated into too many groups, the resulting grouping structure would fail to explain the kinematic connectivity. Therefore, we penalize complex grouping structure.
Denote a forest structure by $\mathcal{F}$, and $F_{d}\in\widetilde{\mathcal{G}}$ as the $d$-th layer’s grouping structure in $\mathcal{F}$. We develop an iterative GIC criterion, with the layer-wise GIC value defined as
$\mbox{GIC}(F_{d})=N^{-1}\sum\limits_{n=1}^{N}(y^{[d+1]}_{n})^{2}+\tau_{d}|F_{d}|,$
where $|F_{d}|$ denotes the number of groups in $F_{d}$, and $y^{[d+1]}_{n}$’s are the ($d+1$)-th layer’s response residuals in the filt-fPLS procedure based on the grouping structure $F_{d}$. We select $F_{d}$ to minimize $\mbox{GIC}(F_{d})$. Here, $\{\tau_{d}\colon d\geq 1\}$ is a non-increasing sequence with respect to $d$, thus the number of groups tends to increase as $d$ increases. The number of tuning parameters $\{\tau_{d}\colon d\geq 1\}$ increases with $d$, making tuning parameter selection increasingly complicated when the total number of layers is large. To solve this issue, we propose adopting a parametric form for $\tau_{d}$ such as $\tau_{\theta}(d)=\rho\gamma^{-d+1}$, where $\theta=(\rho,\gamma)$. The parametric form significantly reduces the computational burden, as it requires selecting only $\theta$ rather than individually determining each $\tau_{d}$. Clearly, denote $\widetilde{\mathcal{F}}$ as the identified forest structure, we have the relationship $\widetilde{\mathcal{F}}\subseteqq\widetilde{\mathcal{G}}\subseteqq\mathcal{P}$.
We now turn to the stopping criterion for the filtrated grouping procedure. In principle, the filtration should continue until the extracted homogeneous components no longer significantly improve the model’s explanatory power. Thus, we propose the following ending criterion: we stop filtration at the $d$-th layer if $\sum_{n}\{(y_{n}^{[d]})^{2}-(y_{n}^{[d+1]})^{2}\}/\sum_{n}(y_{n}^{[d]})^{2}<e$, where $e>0$ is a small value. In addition, if $\sum_{n}\|X^{[d+1]}_{jn}\|^{2}$ becomes sufficiently small, the $j$-th covariate is removed from filtration starting at the $(d+1)$-th layer to reduce computational burden.
3 Simulation results
3.1 General Setting
In this section, we investigate the finite-sample performance of filtrated grouping. It is important to emphasize that, although our simulation is based on a pre-specified forest structure, the objective is not to exactly recover the structure. Rather, the goal is to identify a suitable structure that highlights the numerical advantages of the filtrated grouping framework. The oracle structure is never known in practice, and the forest-structured groups are difficult to fully recover. Nevertheless, our simulation results demonstrate that even when the oracle structure is not fully recovered, the selected forest structure still effectively (even superiorly) captures the underlying homogeneity.
We simulate $N$ samples for the following multiple functional regression model with 10 covariates $\{X_{jn}(t)\colon j=1,\ldots,10\}$, $y_{n}=\sum_{j=1}^{10}\langle X_{jn},\beta_{j}\rangle+\epsilon_{n},$ where $\epsilon_{n}\overset{i.i,d.}{\sim}\mathcal{N}(0,\sigma^{2})$ and $\sigma=0.3,0.5,0.7,1.0$. The covariate and coefficient functions are simulated by the following basis expansion,
$X_{jn}(t)=\sum_{d=1}^{D}\xi_{jn,d}B_{d}(t),\ \beta_{j}(t)=\sum_{d=1}^{D}b_{jd}B_{d}(t),$
where $\{B_{d}(t)\colon d\geq 1\}$ are the Fourier basis functions. The scores $\xi_{jn,d}$ follow the standard normal distribution and are independent across $n$, $j$ and $d$. To facilitate the detection of coefficient homogeneity among the tail scores, we do not impose any decay on the covariate score variance. Here we set $D=9$, and the coefficient scores $\{b_{jd}\colon j=1,\ldots,10,\ d=1,\ldots,D\}$ are displayed in Figure 3.
Figure 3: Coefficient scores $\{b_{jd}\colon d=1,\ldots,D\}$ of the 10 coefficients.
Note that the coefficient scores exhibit partial homogeneity. Specifically, the first two scores are set to 2 for all $j=1,\ldots,p$, reflecting the global homogeneity shared by all covariates. The third and fourth scores are identical within certain subsets of covariates: they are the same for covariates 1 – 3, covariates 4 – 8, and covariates 9 – 10. The last five scores are all different across $j$, reflecting the idiosyncratic effects of different covariates. The $\ell^{2}$-distance between different pairs of coefficient score vectors are displayed in Table 1.
Table 1: $\sum_{d\geq 1}(b_{id}-b_{jd})^{2}$ for $1\leq i<j\leq 10$.
$i$ $j$ 2 3 4 5 6 7 8 9 10
1 0.166 0.337 0.638 0.720 0.699 0.993 0.919 1.523 1.495
2 $\star$ 0.283 0.742 0.754 0.562 0.804 0.748 1.616 1.582
3 $\star$ $\star$ 0.701 0.635 0.704 0.627 0.607 1.382 1.408
4 $\star$ $\star$ $\star$ 0.034 0.224 0.298 0.260 0.387 0.400
5 $\star$ $\star$ $\star$ $\star$ 0.222 0.174 0.158 0.429 0.455
6 $\star$ $\star$ $\star$ $\star$ $\star$ 0.357 0.301 0.412 0.386
7 $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ 0.004 0.580 0.615
8 $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ 0.532 0.560
9 $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ $\star$ 0.005
In this setting, the homogeneity among the 10 coefficient functions cannot be captured by a fixed grouping structure, as the homogeneity across different dimensions is characterized by varying grouping structures. The homogeneity is represented by the forest structure shown in Figure 4. However, this structure may not be optimal. Small differences between $\beta_{4}(t)$ and $\beta_{5}(t)$, $\beta_{7}(t)$ and $\beta_{8}(t)$, as well as $\beta_{9}(t)$ and $\beta_{10}(t)$, could cause these pairs to be merged across more layers in the identified forest structure.
Figure 4: The oracle forest structure underlying coefficient homogeneity.
3.2 Forest Structure Identification
We set $N=400$ and $\sigma=0.5$. The identified forest structure is given in Table 2. The associated filt-fPLS basis functions are displayed in Figure 5, revealing the common variation patterns which determine the associated homogeneous components. For instance, global homogeneity is positively driven by the collective influence of covariates during the intervals 0–0.4 and 0.95–1 of the domain. It is noted that only in the first layer all covariates are clustered into the same group. This global homogeneous component is driven by the first two layers of the oracle structure. In the second layer, the 4th to 10th covariates are clustered together due to their closer values of $b_{jd}$. Specifically, as $d=3,4$, $b_{jd}=1.5$ for $j=4,\ldots,8$ and $b_{jd}=1.8$ for $j=9,10$, and these values are relatively closer to each other compared to $b_{jd}=1$ for $j=1,2,3$. In the 3rd and 4th layers, the closeness between $\beta_{4}(t)$ and $\beta_{5}(t)$, $\beta_{7}(t)$ and $\beta_{8}(t)$, as well as $\beta_{9}(t)$ and $\beta_{10}(t)$, leads to these pairs of covariates being clustered in the same group. In the final layer, the grouping structure captures the individual contributions of each covariate. Overall, the identified forest structure closely aligns with the oracle forest structure and captures the underlying homogeneity more efficiently, as it comprises only six layers and delivers substantially better predictive performance (see Section 3.3).
Table 2: The identified group indices $c_{j,d}$ of different covariates.
$d$ $j$ 1 2 3 4 5 6 7 8 9 10
1 1 1 1 1 1 1 1 1 1 1
2 1 1 1 2 2 2 2 2 2 2
3 1 2 3 4 4 5 6 6 7 7
4 1 2 3 4 4 5 6 6 7 7
5 1 2 3 4 5 6 7 7 8 8
6 1 2 3 4 5 6 7 8 9 10
Figure 5: The estimated fPLS basis functions for all the 6 layers.
3.3 Model Performance and Comparison
To illustrate the superiority of the new filtrated modeling procedure, we compare the prediction performance of the following methods: 1). MFR based on the identified forest structure (Identified); 2). MFR based on the oracle forest structure (Oracle); 3). Grouped multiple function regression (Grouped); 4).  MFR based on the naive grouping structure (NPLS), where no covariates are clustered together in all layers. We test this naive structure to show the necessity of grouping; 5). Ordinary multiple functional linear regression (MFLR).
We repeat 300 simulation runs for each simulation setup. In each run, we generate either 100 or 300 samples for the training set to train the model, and 200 additional samples in the test set used for prediction. The prediction mean squared error (MSE), calculated as $\text{MSE}=\sum_{n=1}^{200}(\hat{y}_{n}-y_{n})^{2}/200,$ is computed for each simulation run. As a result, a total of 300 prediction MSE values are obtained. The box plots of prediction MSEs are presented in Figure 6. They clearly demonstrate that the identified grouping structure outperforms the oracle structure in capturing the relationship between covariates and the response. Its substantial superiority over the naive structure underscores the importance of grouping. As a baseline, the ordinary multiple functional regression model performs the worst. Although the grouped model performs reasonably well, it still falls short compared to the filtrated grouped model. This is because the homogeneity among functional covariates is multi-resolutional, making a fixed grouping structure across different layers inappropriate. Interestingly, although the data are generated based on the oracle structure, the identified forest structure yields significantly better performance. This improvement arises because the data-driven structure more effectively captures the underlying mechanisms of the data. In particular, covariates exhibiting subtle nuances may be better modeled within the same group, thereby rendering the oracle structure suboptimal.
Figure 6: Prediction MSEs of different methods.
4 Filtrated Kinematic Connectivity Analysis
Motivated by its relevance to communicating chronic lower-limb risk associated with aging, we apply the proposed filtrated grouping framework to investigate the relationship between chronological age and angular kinematics using the data collected from the experiment conducted by Fukuchi et al. [14]. Because the sample size is limited, we average the angular kinematic trajectories from the left and right sides. This reduces model complexity while retaining focus on kinematic connectivity. Consequently, there are 15 kinematic trajectories per gait cycle, each serving as a functional covariate, including pelvic obliquity, pelvic rotation, pelvic tilt, hip add/abduction, hip int/external rotation, hip flexion/extension, knee add/abduction, knee int/external rotation, knee flx/extension, ankle inv/eversion, ankle add/abduction, ankle dorsi/plantarflexion, foot inv/eversion, foot int/external rotation and foot DF/plantarflexion angles (see Figure 7). To improve model interpretability, we centralize the log-transformed age and use it as the response $y_{n}$, and standardize the angular kinematic curves as follows,
$\widetilde{X}_{jn}(t)=\frac{X_{jn}(t)-\overline{X}_{j}(t)}{\sqrt{{N^{-1}}\sum\limits_{n=1}^{N}\|X_{jn}(t)-\overline{X}_{j}(t)\|^{2}}},\ j=1,\ldots,15,$
where $\overline{X}_{j}(t)=N^{-1}\sum_{n=1}^{N}X_{jn}(t)$. This preprocessing ensures that group formation reflects the homogeneity of structural and temporal movement patterns, rather than being biased by differences in numerical angular range.
Figure 7: Angular kinematic trajectories averaged across the left and right sides, with curve colors indicating walking speed as shown in Figure 1.
Here, we focus on the comfortable walking speed, specifically within the range of 100–115% of normal walking speed. We employ a Monte Carlo cross-validation approach (see Xu and Liang [37]) to select the tuning parameters in the GIC criterion. Specifically, the samples are randomly split into a training set (80%) and a test set (20%). The model trained on the training set is used to predict the responses in the test set. We repeat this procedure 1000 times to obtain the average prediction MSE. The tuning parameters $\theta$ associated with the minimal prediction MSE is then used for forest structure identification, resulting in the forest structure displayed in Figure 8. The group indices across different filtration layers are shown below:
  • $\widetilde{F}_{1}$.
    $\mathcal{K}_{1,1}$: covariate 1, 4; $\mathcal{K}_{1,2}$: covariate $2,3,5,\ldots,15$;
  • $\widetilde{F}_{2}$.
    $\mathcal{K}_{2,1}$: covariate 1; $\mathcal{K}_{2,2}$: covariate $2,3,5,6,8,\ldots,15$; $\mathcal{K}_{2,3}$: covariate 4; $\mathcal{K}_{2,4}$: covariate 7;
  • $\widetilde{F}_{3}$.
    $\mathcal{K}_{3,i}$: covariate $i$, for $i=1,\ldots,p$;
  • $\widetilde{F}_{4}$.
    $\mathcal{K}_{4,1}$: covariate 2; $\mathcal{K}_{4,2}$: covariate 5; $\mathcal{K}_{4,3}$: covariate 10; $\mathcal{K}_{4,4}$: covariate 11; $\mathcal{K}_{4,5}$: covariate 13; $\mathcal{K}_{4,6}$: covariate 14; $\mathcal{K}_{4,7}$: covariate 15;
  • $\widetilde{F}_{5}$.
    $\mathcal{K}_{5,1}$: covariate 11.
Figure 8: The identified forest structure under 100–115% normal walking speed (the numbers in the circles represent the covariate indices). There are five layers in total. Most covariates are removed from Layers 4 and 5, as the extracted homogeneous components already account for a substantial proportion of their variation.
Define the partial sum of squares (PSS) of the homogeneous components of layer $d$ group $i$ as $\mbox{RSS}_{d,i}$, where $\mbox{RSS}_{d,i}$ represents the residual sum of squares of the reduced model with the corresponding homogeneous component omitted from the full model (all the identified homogeneous components are involved). We use the partial sum of squares to evaluate the importance of the homogeneous components, and a larger value of PSS indicates greater importance. We display the 95% confidence interval of the coefficient scores $\{a_{d,i}\colon i,d\geq 1\}$ and the partial sum of squares of all the identified homogeneous components in Figure 9. It should be noted that all coefficient scores in the first two layers are significantly non‑zero, confirming our hypothesis that kinematic connectivity imparts a consistent age-associated pattern across different kinematic trajectories.
Figure 9: The confidence intervals for coefficient scores and the PSS values. Different layers are separated by blue dashed lines. LdGi represents the $i$-th group in the $d$-th layer, same as $\mathcal{K}_{d,i}$. The insignificant homogeneous components are marked in red.
The homogeneity of the age-kinematics association is captured by the homogeneous components in $\mathcal{K}_{1,1}$, $\mathcal{K}_{1,2}$, and $\mathcal{K}_{2,2}$. In the first hierarchical layer, covariate 1 and 4, representing pelvic obliquity and hip adduction/abduction in the frontal plane, are separated from the other angular kinematic trajectories. Their associated filt‑fPLS basis function (see Figure 10) exhibit pronounced divergence from the one of the other group in the same layer, underscoring that pelvic and hip frontal-plane dynamics exhibit distinct association with age. This reflects the corresponding collective neuromuscular and postural adaptations: during aging, decreased mediolateral balance control and altered weight distribution drive compensatory adjustments in pelvis and hip frontal-plane motion. The corresponding PSS value is considerably higher than those of other homogeneous components, indicating that pelvic and hip frontal-plane motion coordinate as the primary kinematic feature for assessing joint effective age. Meanwhile, we find that the idiosyncratic kinematic features (pertaining to the homogeneous components identified for groups containing a single covariate) of knee internal/external rotation, knee flexion/extension, and ankle dorsiflexion/plantarflexion are non-significant in their associations with age. Their associations with age are adequately captured in the first two filtration layers through the homogeneous components shared with other kinematic trajectories, and their idiosyncratic components do not distinguish gait characteristics between younger and older individuals. Interestingly, the PSS value of the homogenous components in $\mathcal{K}_{3,11}$, corresponding to the idiosyncratic feature of ankle adduction/abduction dynamics exhibits a pronounced spike. Additionally, the three homogeneous components in $\mathcal{K}_{3,11}$, $\mathcal{K}_{4,4}$, and $\mathcal{K}_{5,1}$, which also correspond to the idiosyncratic features of ankle adduction/abduction, all show significant effects. These findings suggest that ankle adduction/abduction kinematics also serve as a particularly informative biomarker for evaluating effective age, and its individual adaption to aging is non-negligible.
Here we summarize the key findings in the filtrated grouping structure:
  • Different angular kinematic trajectories significantly exhibit homogeneity in their age-related association patterns.
  • Pelvic obliquity and hip adduction/abduction kinematics coalesce into a distinct homogeneous component, coordinately serving as the primary indicator of lower-limb joint aging and revealing the age-associated adaptations in pelvic and hip frontal-plane motions linked to mediolateral balance control.
  • Other kinematic trajectories show more homogeneous adaptive patterns, pointing toward system‑wide biomechanical changes associated with aging in gait. Idiosyncratic features of different angular kinematic trajectories are also pronouncedly associated with aging, except for knee internal/external rotation, knee flexion/extension, and ankle dorsiflexion/plantarflexion.
The filt-fPLS basis functions $\{\psi_{d,i}(t)\colon d\geq 1,\ i=1,\ldots,m_{d}\}$ are displayed in Figure 10, and provide interpretable insights into how the homogeneous components relate to age over the gait cycle. Figures 9 and 10 provide a clear interpretation of how the corresponding kinematic trajectories are associated with age. For example, in the first layer, the homogeneous component in Group 1 is primarily associated with pelvis obliquity and hip adduction/abduction. It exhibits a positive relationship with these angular kinematic trajectories during 40–70% of the gait cycle and a negative relationship during the remaining phases, forming a bell-shaped pattern. Combining the result that the associated coefficient score is significantly positive, we conclude that the 40–70% period of the aggregated kinematic trajectories is positively associated with age, whereas the remaining phases exhibit a negative association. In contrast, the homogeneous component in Group 2 are positively driven by all other angular kinematic trajectories throughout the entire gait cycle and exhibit positive association with the response. The corresponding filt-fPLS basis is approximately a constant function, capturing the mean level of the aggregated kinematic trajectories.
Figure 10: The partial least square basis functions in the first two layers.
We benchmark the performance of the new methods against two established alternatives: the ordinary functional regression model (1-1) and the grouped functional regression model (2-3), through comparing the prediction error for age. Prediction accuracy is evaluated using Monte Carlo cross-validation. The participants were required to walk on treadmill under different gait speeds, and we categorize them into four groups: 40–55% (S1), 70–85% (S2), 100–115% (S3), and 130–145% (S4) of the normal walking speed. Then we establish multiple functional linear models under different gait speed groups with response $y^{(s)}_{n}$ and covariates $\widetilde{X}^{(s)}_{jn}(t)$, where the superscript $(s)$ indicates the gait speed levels, $y^{(s)}_{n}=\sum_{j\geq 1}\langle\widetilde{X}^{(s)}_{jn},\beta^{(s)}_{j}\rangle+\epsilon_{n}^{(s)}.$ We repeat the Monte-Carlo sampling 1000 times, and the resulting mean squared errors (MSEs) are displayed in Figure 11. The filtrated model achieves optimal predictive performance at 70–85% of normal walking speed, indicating that assessments conducted at a slightly reduced pace relative to an individual’s comfortable walking speed yield the most precise effective age assessment.
Across all gait speed levels, the filtrated model consistently delivers the lowest and most robust MSEs. The ordinary model (1-1) underperforms due to its high parameter complexity, which leads to increased assessment uncertainty. The grouped model (2-3) demonstrates modest improvements over the ordinary model; however, it remains inferior to the filtrated model. This shortfall stems from its failure to accommodate heterogeneous, idiosyncratic associations between individual covariates and response. By grouping covariates consistently over all dimensions, the grouped model masks their distinct contributions. In contrast, the filtrated model’s hierarchical grouping structure effectively distinguishes shared versus covariate-specific effects, allowing for much more nuanced modeling of age-kinematics associations and significantly enhancing both interpretability and evaluation performance.
Figure 11: Box-plots of MSEs of different methods. The variance of the log-transformed age under the four gait speed levels is 0.19, 0.19, 0.19, and 0.17, respectively.
5 Discussion
We propose a novel filtrated grouping procedure for multiple functional regression models with a scalar response and multiple functional covariates, addressing the challenges posed by lower-limb joint effective age evaluation. By introducing a forest-structured grouped model and a data-driven filtrated partial least squares estimation procedure, our approach effectively captures both homogeneous and idiosyncratic effects of covariates from a multi-resolution perspective, providing systemic insights into kinematic connectivity in relation to age. The filtrated grouping structures are inferred from data, rather than being imposed a priori, making the grouping structures effectively reflect the data mechanism. We apply the filtrated grouping framework to investigate the relationship between chronological age and joint motion patterns in healthy subjects, and identify several key factors that reflect joint aging mechanisms both collectively and individually across different kinematic trajectories. Our results demonstrate the potential of this framework to uncover biomechanical adaptations associated with joint aging, supporting the evaluation of lower-limb effective age and facilitating informed decision-making in clinical and research settings. The proposed approach surpasses existing methods in both interpretability and predictive accuracy, and provides a stratified modeling framework to account for complex covariate connectivity.
Since the filtrated grouping structure is multi-resolution, we should not focus on any single layer but instead consider the overall architecture of grouping. If two angular kinematic trajectories are highly connected, they tend to cluster together across many filtration layers. This insight is useful because persistent grouping across multiple layers signals truly meaningful biomechanical relationships, rather than artifacts of a single grouping structure. In other words, kinematic trajectories that remain clustered consistently across multiple layers of filtration robustly unravels the strong connections.
There are some limitations of our proposed approach. Since the grouping structure follows a forest structure, and becomes increasingly complicated when the dimension increases, it is difficult to provide comprehensive theoretical analysis. In addition, since coefficient homogeneity is sensitive to covariate scaling, the grouping structure can vary when the covariates are scaled differently. This limitation is inherent in any grouping algorithm that relies on coefficient equality. Although covariate standardization can help stabilize group identification, it does not fully eliminate this sensitivity. In contrast, Jiao and Chan [17] developed a scaling-invariant grouping pursuit algorithm based on coefficient shape homogeneity instead of coefficient equality, termed coefficient shape alignment, in which the grouping structure is invariant to covariate scaling without requiring any data preprocessing. We will integrate our filtration-based analysis with coefficient shape alignment in future work to resolve this issue.
This study marks the first effort to integrate filtration techniques into multiple functional regression, and our results underscore the substantial potential and impact of this approach. This novel analytical framework opens up a number of promising directions for future research. For example: How does multi-resolution kinematic connectivity vary across different gait speeds? Are there significant differences in this connectivity structure between normotypical and pathological gait? By comparing the filtrated grouping structures obtained with the same tuning parameters across different samples, we obtain a more comprehensive and insightful comparison. A similar idea has also been applied in network analysis (see Lee et al. [22]) to obtain stable network comparison across different filtration scales. The framework is inherently generalizable and can be extended to other applications involving multivariate functional data. Future work will broaden the filtration paradigm to other functional data methodologies, such as functional factor models and multivariate functional data classification.